OpenStudy (anonymous):

1- Is root mean square just the average amplitude of a wave 2- Why isn't the RMS of a sine wave 2/pi?

OpenStudy (zzr0ck3r):

might have more luck in the physics chan

OpenStudy (anonymous):

Yes, but I already have a running question there.

OpenStudy (zzr0ck3r):

ahh, word well I was just letting you know just in case you didn't know. I'm sure someone here can help eventually...

OpenStudy (anonymous):

I kinda forgotten this topics.. but I remembered that V(RMS) is the average amplitude of the wave

OpenStudy (anonymous):

VRMS = Root 2 / V peak i think

OpenStudy (anonymous):

 Surely it should be\[\frac{\int\limits_{0}^{\pi}Asinx dx}{\pi}\], but http://en.wikipedia.org/wiki/Root_mean_square#RMS_of_common_waveforms disagrees. My method works out to 2A/pi, not A/sqrt2